# Applying a STEM Ways of Thinking Framework for Student-generated Engineering Design-based Physics Problems

### v1

@@ 7 Seeking Qualitative Rigor through Reflexivity
Given that metacognitive reflection is central to our framework, we saw value in turning that lens on our own work. We were guided by the six-way guideline offered by Braund et al. in preparing this reflexivity statement [148]. In this context, Markham shares a precious insight:
@@
One key question we continuously grappled with throughout the study was our writing style. Gilgun advocates for a more evocative and personal style in qualitative research writing. Perhaps in future work, we will revisit and more fully engage with her call to explore the “marvelous range of choices” [153, p.261] qualitative researchers have.
@@ 8 Conclusion
At the beginning of this study, we outlined three research questions (RQs) aimed at applying our WoT4EDP framework to student data. In this section, we summarize our findings for each of our three research questions. While qualitative findings are not intended for generalization [10, p.4], our goal is to explore whether the insights from a small group can shed light on potential issues within a larger group in our specific context. We sught answers to our research questions with this objective in mind. In Section 8.1 we review RQ1; in Section 8.2 we review RQ2; and in Section 8.2 we review RQ3.
8.1 Insights from RQ1: Key Contributions
In Section 5.1, seeking answers for our first research question RQ1, we first examined how students framed their problem statements—a crucial step in design-based thinking. A key insight from our analysis is that students benefit from additional scaffolding to grasp terms like stakeholders, criteria, constraints, and metrics. Since these terms are rarely used in physics classrooms, educators may face challenges in integrating ED-based learning. We hope our work resonates with those seeking to bridge this gap. Our coding sheet in Appendix A can be refined and adapted by educators to develop rubrics for similar activities.
@@ 8.1 Insights from RQ1: Key Contributions
8.2 Insights from RQ2: Key Contributions
@@
(i) DBT - Summary
@@
Regarding DBT (see Section 5.2), one area of concern was how students engaged in iterations. In some cases, iterations were overly simplistic, and some teams engaged in only cursory iterations. Although we did not explicitly instruct students to elaborate on the physical dimensions of their design or the mechanics of their contraption, most groups naturally engaged with these aspects, suggesting an intuitive awareness. However, a more intentional focus on these considerations could further enhance their designs. Encouraging students to conduct feasibility studies could also be beneficial, though this would require explicit instruction on what feasibility entails. Having students develop physical models would be ideal, but this is often impractical in large-enrollment courses. An alternative cost effective approach could be to incorporate process-based problems that involve simulations or open-access data analysis.
(ii) SBT - Summary
@@
Regarding SBT (see Section 5.3), very few teams applied the angular momentum principle, likely due to the nature of their chosen problem or the fact that this topic is introduced later in the semester—after students had already defined their problem in Week 10. However, given that this is a physics course, not surprisingly, all groups engaged with physics concepts and principles. Here again, our rubric (see Appendix D), though fairly basic, appears to have provided some structure. That said, more guidance is needed on the depth of application of these principles. There were scattered but notable instances of misconceptions or inaccuracies in students’ work, which is a crucial issue to address in a physics course. While using the Jupyter Notebook has clear advantages, it also results in students performing calculations without paying attention to numerical precision and units—an oversight that could have serious implications in real-world scenarios. Some teams incorrectly typed their equations, leading to incorrect results, yet they did not seem to recognize the errors. Notably, a few teams wrote their calculations by hand, and while their equations were correctly formulated, they still neglected significant digits. Interestingly, their handwritten work generally avoided excessive numerical precision, unlike those of teams who had Jupyter Notebook outputs. Incorporating elements of the CER (Claim-Evidence-Reasoning) framework [154, 34, 155, 156] could help strengthen the design-science connection by encouraging students to articulate and justify their scientific reasoning more explicitly.
(iii) MBT - Summary
@@
Regarding MBT (see Section 5.4), most teams were comfortable using algebraic equations, though occasional errors were observed. Only one team employed brute-force calculus, while a few others used it indirectly through the momentum update principle. Students also appeared comfortable generating graphs and tables using Python, with some teams opting to draw graphs by hand. Additionally, we noted instances of qualitative reasoning in their discussions. Notably, compared to our earlier study [9], students engaged in more detailed discussions of mathematical ideas. It appears that explicitly prompting students to discuss relevant physics principles may have indirectly encouraged them to integrate mathematical reasoning using physics formulas and principles.
@@
8.3 Insights from RQ3: Key Contributions
@@
Notably, in a recent study, Gambrell and Brewe highlight that in the context of CT in physics, “there is still a lot of research to be done in many of the specific physics courses” [157, p.5]. We hope we have meaningfully contributed to the discussion on computational thinking.
@@ 9 Implications
While our primary audience includes researchers and educators, we believe this work also offers students valuable examples and insights into how their peers think. Studies highlight the potential of peer learning [158] to “foster fertile instructional dialogues among peers” [159, p.4], and we hope student readers will find meaning in the thought processes of their unseen peers.
@@
All models are wrong, but some are useful. [160, p.202]
@@ 10 Future Work
Another potential avenue for exploration is how students engage in argumentation. Following Slavit et al. [13], we are considering applying Toulmin’s Claim–Evidence–Reasoning (C–E–R) framework [154] to analyze how students construct and justify their reasoning.
@@ 11 Acknowledgments
Table 7: Overview of deductive coding summary based on the six CT facets of Shute et al. Each of the 14 teams was coded according to this framework. This table provides a broad representative summary of the teams’ work. See Section (iii) for details.

### v2

@@ 7 Seeking Qualitative Rigor through Reflexivity
Given that metacognitive reflection is central to our framework, we saw value in turning that lens on our own work. While some authors [21, 30] recommend including a positionality statement in qualitative writing, we follow Braund et al. [158] in referring to this section as a reflexivity statement. Their six-way guideline not only shaped the preparation of this statement but also informed our research process more broadly. Since reflexivity is an ongoing requirement throughout the research, an apparent overlap with earlier discussions (see Section 3.3 and Section 4.4) is inevitable. In many ways, reflexivity extends beyond this section, as our methodological discussions also engage with and reinforce reflective considerations throughout the paper.
In this context, Markham shares a precious insight:
@@
This perspective aligns with Barth-Cohen et al. encourage qualitative physics education researchers to include positionality statements, highlighting how their experiences, research roles, and personal knowledge of the study’s topic or domain may have influenced the work [21, p.10].
@@
One key question we continuously grappled with throughout the study was our writing style. Gilgun advocates for a more evocative and personal style in qualitative research writing. Perhaps in future work, we will revisit and more fully engage with her call to explore the “marvelous range of choices” [163, p.261] qualitative researchers have.
@@ 8 Conclusion
At the study’s outset, we posed three research questions (RQs) to apply our WoT4EDP framework to student data. Here, we summarize our findings for each. While qualitative findings are not intended for generalization [10, p.4], we aimed to explore whether insights from a small group could illuminate broader issues in our context. Our analysis was guided by this objective. In Section 8.1 we review RQ1; in Section 8.2 we review RQ2; and in Section 8.2 we review RQ3.
8.1 What do students’ ED problem statements reveal?
In Section 5.1, seeking answers for our first research question RQ1, we first examined how students framed their problem statements—a crucial step in design-based thinking.
Key themes include diverse problem selection, creative framing, stakeholder considerations, and real-world relevance. Teams incorporated numerical metrics, constraints, and assumptions, with some showing an entrepreneurial mindset. All problems had connections to physics and mathematics.
Students would benefit from additional scaffolding to grasp terms like stakeholders, criteria, constraints, and metrics. Since these terms are rarely used in physics classrooms, educators may face challenges in integrating ED-based learning. We hope our work resonates with those seeking to bridge this gap. Our coding sheet in Appendix A can be refined and adapted by educators to develop rubrics for similar activities.
@@ 8.1 What do students’ ED problem statements reveal?
8.2 How do students engage with design, science, mathematics, and metacognitive reflection in their ED problem?
@@
(i) Design-Based Thinking (DBT) – Summary
Regarding DBT (see Section 5.2), key themes include students’ awareness of design criteria, constraints, and numerical details, with varying levels of iteration and refinement. Most teams cited references, but safety and economic considerations were less emphasized. While some teams iterated meaningfully, others showed minimal engagement. Trade-offs were not well-articulated, suggesting a need for better guidance. Overall, students demonstrated design thinking aligned with Crilly’s definition [135], balancing creativity and analytical problem-solving.
@@
One area of concern was how students engaged in iterations. In some cases, iterations were overly simplistic, and some teams engaged in only cursory iterations. Although we did not explicitly instruct students to elaborate on the physical dimensions of their design or the mechanics of their contraption, most teams naturally engaged with these aspects, suggesting an intuitive awareness. However, a more intentional focus on these considerations could further enhance their designs. Encouraging students to conduct feasibility studies could also be beneficial, though this would require explicit instruction on what feasibility entails. Having students develop physical models would be ideal, but this is often impractical in large-enrollment courses. An alternative cost effective approach could be to incorporate process-based problems that involve simulations or open-access data analysis.
(ii) Science-Based Thinking (SBT) – Summary
Regarding SBT (see Section 5.3), key themes include the use of physics principles and concepts, mention of assumptions and approximations, identification of system and surroundings, reporting results, instances of errors, inaccuracies, and oversights.
@@
Very few teams applied the angular momentum principle, likely due to the nature of their chosen problem or the fact that this topic is introduced later in the semester—after students had already defined their problem in Week 10. However, given that this is a physics course, not surprisingly, all teams engaged with physics concepts and principles. Here again, our rubric (see Appendix D), though fairly basic, appears to have provided some structure. That said, more guidance is needed on the depth of application of these principles. There were scattered but notable instances of misconceptions or inaccuracies in students’ work, which is a crucial issue to address in a physics course. While using the Jupyter Notebook has clear advantages, it also results in students performing calculations without paying attention to numerical precision and units—an oversight that could have serious implications in real-world scenarios. Some teams incorrectly typed their equations, leading to incorrect results, yet they did not seem to recognize the errors. Notably, a few teams wrote their calculations by hand, and while their equations were correctly formulated, they still neglected significant digits. Interestingly, their handwritten work generally avoided excessive numerical precision, unlike those of teams who had Jupyter Notebook outputs. Incorporating elements of the CER (Claim-Evidence-Reasoning) framework [164, 34, 165, 166] could help strengthen the design-science connection by encouraging students to articulate and justify their scientific reasoning more explicitly.
(iii) Mathematics-Based Thinking (MBT) – Summary
Regarding MBT (see Section 5.4), key themes include how students performed and presented their calculations, and how they reported their findings.
@@
Most teams were comfortable using algebraic equations, though occasional errors were observed. Only one team employed brute-force calculus, while a few others used it indirectly through the momentum update principle. Students also appeared comfortable generating graphs and tables using Python, with some teams opting to draw graphs by hand. Additionally, we noted instances of qualitative reasoning in their discussions. Notably, compared to our earlier study [9], students engaged in more detailed discussions of mathematical ideas. It appears that explicitly prompting students to discuss relevant physics principles may have indirectly encouraged them to integrate mathematical reasoning using physics formulas and principles.
@@
8.3 In what ways do students demonstrate computational thinking in their ED challenges?
@@
Notably, in a recent study, Gambrell and Brewe highlight that in the context of CT in physics, “there is still a lot of research to be done in many of the specific physics courses” [167, p.5]. We hope we have meaningfully contributed to the discussion on computational thinking.
@@ 9 Implications
While our primary audience includes researchers and educators, we believe this work also offers students valuable examples and insights into how their peers think. Studies highlight the potential of peer learning [168] to “foster fertile instructional dialogues among peers” [169, p.4], and we hope student readers will find meaning in the thought processes of their unseen peers.
@@
All models are wrong, but some are useful. [170, p.202]
@@ 10 Future Work
Another potential avenue for exploration is how students engage in argumentation. Following Slavit et al. [13], we are considering applying Toulmin’s Claim–Evidence–Reasoning (C–E–R) framework [164, 165] to analyze how students construct and justify their reasoning.
@@ 11 Acknowledgments
Table 7: Deductive coding summary based on the six CT facets of Shute et al. Each of the 14 teams was coded separately according to this framework. This table provides a broad representative summary of all the teams taken as a whole. See Section (iii) for details.